OpenStudy (anonymous):

I WILL FAN AND MEDAL!!!! twelve to thirty−two Select one: A. nine over twenty B. fifteen colon forty C. ten over twenty−five D. 6 to 8

OpenStudy (anonymous):

@crazydave23

ganeshie8 (ganeshie8):

cross multiply

OpenStudy (anonymous):

cross mustiply what?

OpenStudy (anonymous):

@raymonde20

OpenStudy (anonymous):

Fractions.

OpenStudy (welshfella):

or make it into a decimal 32 / 12 = ?

ganeshie8 (ganeshie8):

lets test first option by cross multiplying : is below true ?\[\large \dfrac{12}{32}~~=~~\dfrac{9}{20}\] \[\large 12\times 20~~ =~~ 9\times 32\]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

really ? 12*20 = ? 9*32 = ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh sorry i though did the * number match

ganeshie8 (ganeshie8):

good, so first option is not the answer lets test second option

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

lets test second option by cross multiplying : is below true ?\[\large \dfrac{12}{32}~~=~~\dfrac{15}{40}\] \[\large 12\times 40~~ =~~ 15\times 32\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's B!

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

thx a baunch

ganeshie8 (ganeshie8):

yw!